OpenStudy (anonymous):

Need help, please. Jean and Mark are going to fill a pool with 2 different sized hoses. Jean can fill the pool in 8 hours, while Mark can complete it in 12 hours. Their supervisor thinks that the job will take 10 hours to complete if they work together. Explain each step in solving this equation and determine if the supervisor is correct or not.

OpenStudy (mertsj):

Jean fills 1/8 of the pool per hour. Do you understand that?

OpenStudy (anonymous):

Yes, and then mark it would be 1/12?

OpenStudy (mertsj):

Yes.

OpenStudy (mertsj):

So what part of the pool could Jean fill in 3 hours?

OpenStudy (anonymous):

1/3 ?

OpenStudy (mertsj):

no.

OpenStudy (mertsj):

This is such a key concept. Let's see if you can understand it in a different way. Have you ever had a job of any kind? Maybe babysitting or something?

OpenStudy (anonymous):

I've watched my brothers dog a few times

OpenStudy (mertsj):

How much did you make each day for watching the dog?

OpenStudy (anonymous):

Well, he's still a puppy so I had to watch him usualy from 8 in the morning until 9 when my mom got home

OpenStudy (mertsj):

Did you get paid?

OpenStudy (anonymous):

Usually he would buy me a new game or donuts or something lol

OpenStudy (mertsj):

Well, let's just say you had a babysitting job and you made $5 per hour. Then you would make $5 times 3 in 3 hours. Do you understand that?

OpenStudy (anonymous):

Yes, because if it's 5 dollars an hour and I worked for 3, i would do the 5 x 3 to find out how much I would make?

OpenStudy (mertsj):

Yes. Very good.

OpenStudy (mertsj):

Now if you baby sat for h hours you would make $5 times h or 5h dollars. Do you understand that?

OpenStudy (anonymous):

Yes, because in this situation you don't know how many hours I've worked

OpenStudy (mertsj):

Now let's go back to the pool.

OpenStudy (mertsj):

Jean fills 1/8 pool each hour. So in x hours she would fill 1/8 times x or (1/8)(x) or \[\frac{x}{8}\] pools

OpenStudy (mertsj):

And if you understand that, then tell me how many pools Mark fills in x hours

OpenStudy (anonymous):

If mark fills 1/12 of pool each hour. Then mark would in x hours fill 1/12 times x or (1/12)(x) and it would be \[\frac{ x }{ 12} ?\]

OpenStudy (mertsj):

Perfect

OpenStudy (mertsj):

Now together they fill 1 pool so add what Jean does to what Mark does and set it equal to 1

OpenStudy (anonymous):

so \[\frac{ x }{ 8} + \frac{ x }{ 12 }\] thats how to set it up right?

OpenStudy (mertsj):

Yes and put = 1

OpenStudy (anonymous):

alright so, would I need to find a common denominator?

OpenStudy (mertsj):

Yes. Which is 24. So multiply both sides of the equation by 24

OpenStudy (anonymous):

\[\frac{ x }{ 8 } (24) + \frac{ x }{ 12 } (24) = 1\] so now it's \[\frac{ 24x }{ 24 } + \frac{ 24x}{ 24}\] sorry if I get this totally incorrect

OpenStudy (mertsj):

\[\frac{x}{8}\times \frac{24}{1}= 3x\] because the 8 cancels into the 24

OpenStudy (mertsj):